OpenStudy (anonymous):

Please help, for medal and fan!!!!!!!!!!!!!!!!!! Evaluate the logarithm. log(3)1/27

OpenStudy (anonymous):

\[\log_3{\frac{ 1}{27 }} \]

OpenStudy (mathmale):

Apply this rule:\[\log \frac{ a }{ b }=\log a - \log b\]

OpenStudy (mathmale):

Also use this second rule:\[\log_{a}a^b=b \]